OpenStudy (anonymous):

Ethylenediamine (C2H8N2) is a liquid reagent and the concentrated form of this reagent has a density of 0.899 g/mL. What volume of concentrated ethylenediamine should be diluted with water to prepare 100.0 mL of a 0.150 M solution of ethylenediamine?

OpenStudy (aaronq):

First find the moles of solute you need. \(Molarity_{solution}=\dfrac{n_{solute}}{L_{solution}}\) then find the mass. \(n=\dfrac{m}{Molar\;mass}\) then find the volume using the density \(\rho=\dfrac{m}{V}\)